OpenStudy (anonymous):

24 base 6 multiplied by 5 base 6. (Answer in base 6.)

OpenStudy (welshfella):

I'm struggling to remember number systems my first thought is to convert to decimal first Then do the multiplication and convert decimal back to base 6

OpenStudy (anonymous):

I did not know that was even doable. I will look into that. Thank you.

OpenStudy (welshfella):

24 in base 6 = 2*6^1 + 4*6^0 = 12 + 4 = 16 in base 10 5 in base 6 = 5 in base 10 16 * 5 = 80 base 10

OpenStudy (welshfella):

now to convert this to base 6 is the next step... 80 / 6 = 13 remainder 2 so the last digit is 2

OpenStudy (welshfella):

now we need to make 78 form x * 6^2 + y * 6^1 = 1* 36 + 7 * 6^1 = 78 so the answer is 172 base 6

OpenStudy (welshfella):

- there is probably an easier way of doing this....